OpenStudy (anonymous):

Hi I need some help cause I'm confused. The is problem is if 7/8 liters fills 5/6 bottle of water. How much liters fills the whole bottle?

OpenStudy (mathteacher1729):

Basically it is asking you to solve this equation for b: \[\frac{7}{8}L=\frac{5}{6}b\] The answer will be "some fraction" times L = 1 bottle

OpenStudy (anonymous):

Oh so it is just a simple proportion?

OpenStudy (mathteacher1729):

Yup.

OpenStudy (mathteacher1729):

They try to throw you off with the weird fractions, but ultimately it's just proportion.

OpenStudy (anonymous):

The wording confused me.

OpenStudy (anonymous):

So the answer would be 1 \[1\frac{ 1 }{ 20 }?\]

OpenStudy (mathteacher1729):

Yup. That's it.

OpenStudy (anonymous):

Oh Thank you. I thought that was the answer but wasn't 100% sure